OpenStudy (anonymous):

oxygen used in hospitals and laboratories is often obtained from cylinders containing liquefied oxygen. a cylinder contained 1.00 x 10^2kg of liquid oxygen. the volume(litters) of oxygen that could produced at 1.00atm of pressure at room temperature, 20.0 degree celcius is, (atomic mass of O=16g, R=0.08205 L atm mol^-1 K^-1)

OpenStudy (abmon98):

Whats your question?

OpenStudy (anonymous):

the volume(litters) of oxygen that could produced at 1.00atm of pressure at room temperature, 20.0 degree celcius is @Abmon98

OpenStudy (abmon98):

1*10^5*V=0.08205*(273+20)*(1*10^2*10^3/16)

OpenStudy (anonymous):

1*10^2*10^3/16) this part that 16 is divide by 10^3 only or the whole bracet? @Abmon98

OpenStudy (abmon98):

nope the whole bracket because we need to change the kg into grams to get the number of moles

OpenStudy (anonymous):

the answer u gt was it 1.5025??

OpenStudy (abmon98):

yes exactly

OpenStudy (anonymous):

but this answer is not in the choices given!! A. 75.2X10^2 L B. 75.2X10^3 L C. 0.752 L D. 75.2 L E. 7.52 L @Abmon98

OpenStudy (abhisar):

Use PV = nRT

OpenStudy (matt101):

I get B as my answer. Using PV=nRT and solving for V: P = 1 atm n = 100 000 g / 32 g/mol = 3125 mol R = 0.08206 T = 293 K